OpenStudy (anonymous):

Find the solution of the initial value problem. dy/dx = 9x^2 + 8x; y(3) = 15

OpenStudy (anonymous):

mult both side by dx, integrate, adjust for initial condition: y(3) = 15

OpenStudy (anonymous):

the answer i got was 3x^3+4x^2-129

OpenStudy (anonymous):

it turned out to be wrong

OpenStudy (anonymous):

3(3)^3 +4(3)^2 + c = 15 => 81 +36 + c = 15 => c = 15 - 117 = -112

OpenStudy (anonymous):

you got the wrong constant...check you're arithmetic

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

yeah i am! sorry! i just plugged in the constant and the answer was still wrong.. I'll do it again to see if i get it

OpenStudy (anonymous):

y(x) = 3x^3+4x^2-112 y(3) = 15 dy/dx = 9x^2 + 8x

OpenStudy (anonymous):

so the original functions is 3x^3+4x^2 +c since its the antiderivative Then you have to find C so you plug in the x value into the original function and set it equal to 0 so you get --> 3(3)^3+4(3)^2=15 right..? that how you got c=112 right?

OpenStudy (anonymous):

i put the answer you got and it was wrong as well

OpenStudy (anonymous):

yeah... 3(3)^3+4(3)^2+c=15 => 81+36+c = 15 => 117 + c = 15 => c = -102 sorry, I should have check MY arithmetic too!!!